OpenStudy (anonymous):

Find the fifth term of the sequence which has a first term of 17 and has a common ratio of 3.1 Round your answer to the nearest hundredth if necessary.

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

@campbell_st

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle a_n=a_1\times r^{n-1} }\)

OpenStudy (solomonzelman):

plug in your numbers, that's all.

OpenStudy (solomonzelman):

(note, you need: \(\large\color{slate}{\displaystyle a_5 }\) )

OpenStudy (anonymous):

I'm confused @SolomonZelman